OpenStudy (anonymous):

Find the slope of the curve 2(x+y)^1/3=y at the point (4,4)

OpenStudy (therealmeeeee):

Take the derivative. y^3 - xy^2 = 4 3y^2(dy/dx) - (y^2 + 2xy(dy/dx)) = 0 Isolate dy/dx. 3y^2(dy/dx) - 2xy(dy/dx) = y^2, factor dy/dx (dy/dx)(3y^2 - 2xy) = y^2, divide both sides by (3y^2 - 2xy) dy/dx = y^2/(3y^2 - 2xy) is your derivative equation. Now we need x. Substitute 2 for y in the original equation and solve for x. (2)^3 - x(2)^2 = 4 4x = 8 - 4 4x = 4 x = 1 Now, substitute x and y into the derivative's equation. dy/dx = (2)^2/(3(2)^2 - 2(1)(2)) = 4/(12 - 4) = 1/2 The slope, dy/dx, is 1/2 when y = 2.

OpenStudy (therealmeeeee):

there ya go

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

but this wasnt my problem

OpenStudy (anonymous):

@TheRealMeeeee